Wildhorse Co. began operations on July 1. It uses a perpetual inventory system. During July, the company had the following purchases and sales.

Purchases

Date

Units

Unit Cost

Sales Units

July 1 5 $166
July 6 3
July 11 6 $180
July 14 5
July 21 7 $191
July 27 5
Calculate the average cost per unit at July 1, 6, 11, 14, 21 & 27. (Round answers to 2 decimal places, e.g. $105.50.)

Average cost for each unit

July 1 $

July 6 $

July 11 $

July   14 $

July 21 $

July 27 $

LINK TO TEXT

Determine the ending inventory under a perpetual inventory system using (1) FIFO, (2) moving-average cost, and (3) LIFO. (Round average-cost per unit to 2 decimal places, e.g. 12.52 and final answer to 0 decimal places, e.g. 1,250.)

FIFO

MOVING-AVERAGE

LIFO

The ending inventory under a perpetual inventory system $

$

$

LINK TO TEXT

Which costing method produces the highest ending inventory valuation?

LIFO method Average-cost method FIFO method
Answer #1
Calculate the average cost per unit at July 1, 6, 11, 14, 21 & 27. (Round answers to 2 decimal places, e.g. $105.50.)

Average cost for each unit

July 1 $166
July 6 $166
July 11 (2*166+6*180/8) 176.50
July   14 176.50
July 21 (7*191+3*176.50/10) $186.65
July 27 $186.65

LINK TO TEXT

Determine the ending inventory under a perpetual inventory system using (1) FIFO, (2) moving-average cost, and (3) LIFO. (Round average-cost per unit to 2 decimal places, e.g. 12.52 and final answer to 0 decimal places, e.g. 1,250.)

FIFO

MOVING-AVERAGE

LIFO

The ending inventory under a perpetual inventory system 5*191 = 955

5*186.65 = 933.25

(2*166+1*180+2*191) = 894

Highest inventory method = FIFO